if x-y=20 and x is 30 find y
Answer by mananth(16946)

x-y = 20
x = 30
in place of x put 30
30-y= 20
add -30 to both sides
30-30-y=20-30
-y=-10
divide by -1
-1/1 = -10/1
y=10